Indigo River Technology is considering a project that would last for 2 years. The project would involve an initial investment of 91,000 dollars for new equipment that would be sold for an expected price of 76,000 dollars at the end of the project in 2 years. The equipment would be depreciated to 19,000 dollars over 6 years using straight-line depreciation. In years 1 and 2, relevant annual revenue for the project is expected to be 75,000 dollars per year and relevant annual costs for the project are expected to be 35,000 dollars per year. The tax rate is 50 percent and the cost of capital for the project is 7.41 percent. What is the net present value of the project?
Answer #1

Initial Investment = $ 91000, Equipment Sale Price = $ 76000 and Depreciated Value = $ 19000, Depreciation Tenure = 6 years

Depreciation type: straight line

Annual Depreciation Expense = (91000 - 19000) / 6 = $ 12000

Annual Revenue = $ 75000 and Annual Expense = $ 35000

Project Tenure: 2 years

Annual Revenue = $ 75000

Less: Annual Expense = $ 35000

EBITDA = $ 40000

Less: Depreciation Expense = $ 12000

EBIT = $ 28000

Less: Tax Expense @ 50% = 0.5 x 28000 = $ 14000

Net Income = $ 14000

Add: Depreciation Expense = $ 12000

Operating Cash Flow (OCF) = $ 26000

Cost of Capital = 7.41 %

Book Value of Equipment at the end of Year 2 = 91000 - 12000 - 12000 = $ 67000

After-Tax Salvage Value = 76000 - 0.5 x (76000-67000) = $ $ 71500

Therefore, Project NPV = 26000 x (1/0.0741) x [1-{1/(1.0741)^(2)}] + [71500 / (1.0741)^(2)] - 91000 = $ 17717.69